OpenStudy (anonymous):

Find the total amount if you deposit $500 at a rate of 5% for two years using simple interest. Year 1 Year 2

OpenStudy (anonymous):

year 1 = 500 + 0.05 * 500 = $525 year 2 = 500 + 0.05 * 500 * 2 = 550 It is a common mistake to take interest of $525 at end of year 2. But since it is simple interest hence there should not be any compounding, not even at the end of a year.